# 11footnotetext: co-authors   co-senior authors

### v2

@@ 6 Conclusion and Limitations
Since this version of the DGM is mainly powered by FMs, it is inherently limited by the capabilities of the underlying FM. Hence, an exciting future direction is to extend self-modification beyond just prompts or FM workflows, to include more computationally intensive methods, such as rewriting its own training script to update the FM itself. While this version of the DGM focuses on coding, AI systems are increasingly applied across a wide range of domains (e.g., computer vision, creative writing). Another promising extension is to develop self-improving AI systems capable of enhancing themselves beyond just the coding domain. A key assumption in this work is that coding benchmarks are a good reflection of the agent’s ability to self-improve, since the self-modification task requires the agent to modify its own codebase. However, one could envision an alternative approach that co-evolves the target task distribution (Faldor et al., 2025; Wang et al., 2023c), thereby removing the constraint of self-improvement being tied to a single objective, as in true open-ended processes. Appendix˜I presents additional potential directions for future work. As we continue to explore this powerful technology, we must also keep safety front and center, as discussed in Section˜5.
@@ Appendix B Additional Related Work
Open-Endedness (part 2). Early approaches to open-endedness explored different mechanisms to balance learnability and interestingness. Quality-diversity algorithms sought to illuminate vast solution spaces with diverse, high-performing behaviors (Pugh et al., 2016; Chatzilygeroudis et al., 2021; Mouret & Clune, 2015; Nguyen et al., 2015). Other methods emphasized goal-directed exploration (Ecoffet et al., 2019; 2021; Schaul et al., 2015; Andrychowicz et al., 2017; Eysenbach et al., 2018), intrinsic motivation (Lehman & Stanley, 2011; Oudeyer et al., 2007; Li et al., 2014; Pathak et al., 2017), or learning progress frameworks (Kanitscheider et al., 2021; Gaven et al., 2025; Baranes & Oudeyer, 2013; Colas et al., 2019; 2022b; Jiang et al., 2021; Dennis et al., 2020; Schmidhuber, 2008; 2013; Kompella et al., 2017). More recently, large-scale foundation models (FMs) (Brown et al., 2020; Radford et al., 2019) have emerged as powerful proxies for human notions of interestingness (Zhang et al., 2024b; Faldor et al., 2025; Sancaktar et al., 2025) and effective mutation operators to propose novel solutions in code (Romera-Paredes et al., 2024; Novikov et al., 2025; Lehman et al., 2023; Faldor et al., 2025; Hu et al., 2025). FMs can guide autotelic agents (Colas et al., 2022b; 2023; a), model human preferences for quality and diversity (Bradley et al., 2024; Ding et al., 2024; Wang et al., 2023b; Klissarov et al., 2023; 2024; Samvelyan et al., 2024; Lim et al., 2024; Havrilla et al., 2024a), design reward functions (Wiering & Schmidhuber, 1997; Wang et al., 2023a; Ma et al., 2023; Faldor et al., 2025), create simulated environments (Sudhakaran et al., 2023; Nasir & Togelius, 2023; Aki et al., 2024; Nasir et al., 2024; Bruce et al., 2024; Parker-Holder et al., 2024; Schmidhuber, 2013), drive ever-evolving multi-agent dynamics (Dharna et al., 2024; Zhou et al., 2025), search diverse ambulating robot morphologies (Lehman et al., 2023), and search expansive solution spaces for benchmark or objective optimization (Lange et al., 2024; Zhang et al., 2024b; Faldor et al., 2025; Hu et al., 2025; Lu et al., 2024b; Romera-Paredes et al., 2024; Fernando et al., 2024; Lu et al., 2024a; Khan et al., 2024; Lu et al., 2025; Liu et al., 2024; Novikov et al., 2025).
@@
Program Synthesis. Program synthesis (Alur et al., 2018; Polozov & Gulwani, 2015; Buchi & Landweber, 1990; Gulwani, 2011; Ellis et al., 2021) seeks to generate code meeting external specifications such as input-output examples or logical formulas. Hybrid approaches combine symbolic methods with neural or FM guidance: for instance, Li et al. (2024) uses LLM suggestions to steer symbolic search in SyGuS settings, improving over pure enumeration. Barke et al. (2024) blends LLM completions with a learned surrogate model to guide synthesis in DSLs. Shi et al. (2023) uses neural policies to build higher-order and lambda abstractions during search, outperforming both LLM-only and symbolic baselines on list manipulation tasks. The DGM differs in focusing not just on producing programs for external tasks, but also on agent self-modification, rewriting its own implementation to improve its capacity for future self-improvement.
@@ C.2 Parent Selection
This parent selection favors agents with high performance and fewer existing children, thereby promoting both exploitation and exploration in the open-ended process. All agents in the archive maintain a non-zero probability of being sampled, meaning that any path to improvement is feasible given sufficient compute.
@@ Appendix H Case Study: Solving Hallucination
Appendix I Additional Future Work Directions

### v3

@@ 6 Conclusion and Limitations
Since this version of the DGM is mainly powered by FMs, it is inherently limited by the capabilities of the underlying FM. Hence, an exciting future direction is to extend self-modification beyond just prompts or FM workflows, to include more computationally intensive methods, such as rewriting its own training script to update the FM itself. While this version of the DGM focuses on coding, AI systems are increasingly applied across a wide range of domains (e.g., computer vision, creative writing). Another promising extension is to develop self-improving AI systems capable of enhancing themselves beyond just the coding domain. A key assumption in this work is that coding benchmarks are a good reflection of the agent’s ability to self-improve, since the self-modification task requires the agent to modify its own codebase. However, one could envision an alternative approach that co-evolves the target task distribution (Faldor et al., 2025; Wang et al., 2023c), thereby removing the constraint of self-improvement being tied to a single objective, as in true open-ended processes. Appendix˜J presents additional potential directions for future work. As we continue to explore this powerful technology, we must also keep safety front and center, as discussed in Section˜5.
@@
Ethics Statement
We affirm compliance with the ICLR Code of Ethics. This work studies self-improving AI systems in the limited context of code-editing agents evaluated on standard programming benchmarks. No human subjects were involved and no personally identifiable information (PII) was collected or processed; IRB approval was therefore not required.
Safety and misuse. Self-modifying systems can pose safety risks if allowed to act without constraints or if optimizations inadvertently introduce unsafe behaviors. To mitigate this, all agents in our experiments ran inside isolated sandboxes with strict resource and time limits; agents had limited network access and no ability to modify the host environment. The self-improvement scope was restricted to the agent’s own Python codebase and evaluation harnesses. We maintained a complete, auditable lineage (archive) of code changes and evaluations, enabling rollback and post-hoc analysis. We did not deploy discovered agents in real development environments. Our release plan (code, prompts, and evaluation artifacts) will exclude any components that grant elevated system access and will include default sandboxing, guardrails, and clear documentation of intended use.
Dual-use, downstream impact, and limitations. Stronger autonomous coding agents could be dual-use (e.g., aiding software maintenance, but also potentially facilitating creation of harmful code if misapplied). We believe the research benefits (e.g., advancing methods for controlled, auditable self-improvement and demonstrating practical safeguards) outweigh the risks. Nevertheless, we explicitly discourage security-sensitive or unsandboxed deployment and provide concrete safety recommendations (Section˜5). Our empirical focus on benchmark optimization may not capture all desirable properties (robustness, interpretability, or broader social values). We therefore treat benchmark gains as necessary but insufficient indicators of general AI development, and discuss avenues to integrate other objectives (e.g., safety, reasoning) into the optimization loop.
Data governance, IP, and licensing. We evaluate on SWE-bench Verified and Polyglot, which are composed of open-source repositories and tasks. We complied with dataset licenses and usage terms to the best of our knowledge. We did not introduce or distribute proprietary code. Foundation models (FMs) were accessed via provider APIs under their terms of service; we did not submit sensitive data, nor attempt to circumvent usage policies. Logs released with this work will be scrubbed of API keys and any incidental sensitive strings.
Bias, fairness, and equity. Although our domain is software code rather than human-centered text, FM behavior can still reflect biases (e.g., language or ecosystem preferences) and may unevenly benefit communities whose tooling is better represented in training data. We partially address this by evaluating across multiple languages (Polyglot) and reporting cross-benchmark transfer. Future work should add diagnostics for biased failure modes and include broader, community-driven task sets.
Conflicts of interest and funding. No author has a financial interest in products whose performance is evaluated here. Sponsors and employers did not influence experimental design, analysis, or the decision to publish, beyond providing salary or standard research support. Any external compute or API credits are acknowledged in the appendix.
Reproducibility Statement
We will open-source all code and full agent logs, including the complete archive lineage of self-modifications (diffs, prompts, and configs) as well as the evaluation harness. To support exact replication, we reference the following: algorithmic details and pseudocode (Section˜3, Section˜C.4); parent selection and open-ended exploration settings (Section˜C.2); foundation model choices and hyperparameters (Section˜D.1); benchmark task subsets for SWE-bench and Polyglot (Section˜E.2, Section˜E.3); staged evaluation protocols and scripts (Section˜4.2); implementations and diffs for the best discovered agents (Section˜F.1, Section˜F.2); and compute and cost estimates (Section˜E.1). The released code repository will include environment specifications and scripts to reproduce all results, figures, and tables.
@@ Appendix B Additional Related Work
Open-Endedness (part 2). Early approaches to open-endedness explored different mechanisms to balance learnability and interestingness. Quality-diversity algorithms sought to illuminate vast solution spaces with diverse, high-performing behaviors (Pugh et al., 2016; Chatzilygeroudis et al., 2021; Mouret and Clune, 2015; Nguyen et al., 2015). Other methods emphasized goal-directed exploration (Ecoffet et al., 2019; 2021; Schaul et al., 2015; Andrychowicz et al., 2017; Eysenbach et al., 2018), intrinsic motivation (Lehman and Stanley, 2011; Oudeyer et al., 2007; Li et al., 2014; Pathak et al., 2017), or learning progress frameworks (Kanitscheider et al., 2021; Gaven et al., 2025; Baranes and Oudeyer, 2013; Colas et al., 2019; 2022b; Jiang et al., 2021; Dennis et al., 2020; Schmidhuber, 2008; 2013; Kompella et al., 2017). More recently, large-scale foundation models (FMs) (Brown et al., 2020; Radford et al., 2019) have emerged as powerful proxies for human notions of interestingness (Zhang et al., 2024b; Faldor et al., 2025; Sancaktar et al., 2025) and effective mutation operators to propose novel solutions in code (Romera-Paredes et al., 2024; Novikov et al., 2025; Lehman et al., 2023; Faldor et al., 2025; Hu et al., 2025). FMs can guide autotelic agents (Colas et al., 2022b; 2023; a), model human preferences for quality and diversity (Bradley et al., 2024; Ding et al., 2024; Wang et al., 2023b; Klissarov et al., 2023; 2024; Samvelyan et al., 2024; Lim et al., 2024; Havrilla et al., 2024a), design reward functions (Wiering and Schmidhuber, 1997; Wang et al., 2023a; Ma et al., 2023; Faldor et al., 2025), create simulated environments (Sudhakaran et al., 2023; Nasir and Togelius, 2023; Aki et al., 2024; Nasir et al., 2024; Bruce et al., 2024; Parker-Holder et al., 2024; Schmidhuber, 2013), drive ever-evolving multi-agent dynamics (Dharna et al., 2024; Zhou et al., 2025), search diverse ambulating robot morphologies (Lehman et al., 2023), and search expansive solution spaces for benchmark or objective optimization (Lange et al., 2024; Zhang et al., 2024b; Faldor et al., 2025; Hu et al., 2025; Lu et al., 2024b; Romera-Paredes et al., 2024; Fernando et al., 2024; Lu et al., 2024a; Khan et al., 2024; Lu et al., 2025; Liu et al., 2024; Novikov et al., 2025).
@@
Program Synthesis. Program synthesis (Alur et al., 2018; Polozov and Gulwani, 2015; Buchi and Landweber, 1990; Gulwani, 2011; Ellis et al., 2021) seeks to generate code meeting external specifications such as input-output examples or logical formulas. Hybrid approaches combine symbolic methods with neural or FM guidance: for instance, Li et al. (2024) uses LLM suggestions to steer symbolic search in SyGuS settings, improving over pure enumeration. Barke et al. (2024) blends LLM completions with a learned surrogate model to guide synthesis in DSLs. Shi et al. (2023) uses neural policies to build higher-order and lambda abstractions during search, outperforming both LLM-only and symbolic baselines on list manipulation tasks. The DGM differs in focusing not just on producing programs for external tasks, but also on agent self-modification, rewriting its own implementation to improve its capacity for future self-improvement.
@@ E.1 Cost Estimate
Also, higher-performing agents discovered by the DGM do indeed incur greater inference costs than the initial agent, but cost and performance are not strictly correlated, where some expensive agents underperform cheaper ones.
@@ Appendix H Case Study: Solving Hallucination
Appendix I Additional Safety Discussion
Any advancement that increases the autonomous capabilities of AI systems introduces its own set of safety considerations (Bengio et al., 2024), especially for systems that improve in an open-ended way (Ecoffet et al., 2020; Clune, 2019). Section˜5 discusses these concerns and outlines concrete, actionable steps for mitigating them. We call for much more research into and discussion regarding AI safety, including deep thought and discussion amongst all stakeholders in society on the complicated question of what exactly counts as safe AI. We are confident the work we have done was never unsafe (Section˜5), but scaled up versions of it could be. As with all transformative technologies, the ultimate impact of such AI systems remains deeply uncertain, and good arguments can be made both for the case that it will bring about tremendous good and tremendous harm. These uncertainties highlight the need for sustained, inclusive, and multidisciplinary discussion (not only from current experts but also from a wider and more diverse community) on how to navigate these developments.
Appendix J Additional Future Work Directions
@@ Appendix I Additional Safety Discussion
DGM with Advanced Foundation Models. Recent FMs have advanced dramatically, enabling scaffolds to become simpler on current coding benchmarks (Yang et al., 2024). It is possible that in some settings, like current coding benchmarks, certain engineering efforts in scaffolding might be downplayed by the improvement of the FMs. However, many scaffolding components (advanced tools, parallel workflows, external memory, proxy verification, etc.) still fundamentally can not be internalized by FMs and will be essential for more complex real-world tasks beyond today’s benchmarks. Future work to explore how different components in agents will emerge with different FMs could be a promising direction.
Evolving Generalist Agent. We believe some degree of task-specific adaptation is indeed expected and even desirable, since fundamentally different types of tasks (e.g., in our case, multi-file Python repository edits vs. primarily single-file, multi-language implementations) naturally require distinct scaffolding components. Crucially, this very property highlights a unique advantage of self-improving systems like the DGM: it replaces laborious manual efforts to design specialized agents for diverse tasks with a fully automated evolutionary process. This motivates an exciting future direction of running the DGM on a large, diverse set of tasks to evolve a true generalist agent. Also, currently we only evaluated DGM on two coding benchmarks. While we believe these two benchmarks differ substantially in task structure (multi-file Python repository edits vs. primarily single-file, multi-language implementations), providing strong evidence of generality, additional benchmarks would further strengthen evaluations.